The larger the standard deviation of a distribution, the more heterogeneous the scores in that distribution. Is this statement true?explain.

1
Expert's answer
2022-04-25T08:05:38-0400

It is true. Using "\sigma=\sqrt{\frac{\sum(x-\mu)^2}{n}}" , when the difference between x and the mean is greater, then "\sigma" is greater.